### v1

# Robustness of timelike circular orbit topology against particle spin

Abstract
Based on a detailed study of the motion of spinning test particles within the Mathisson–Papapetrou–Dixon formalism under the Tulczyjew spin-supplementary condition in static, spherically symmetric spacetimes, we investigate the topological properties of timelike circular orbits (TCOs) for such particles. By constructing an auxiliary potential and an associated vector field on the equatorial plane, we compute the topological winding number $W$ for regions between horizons and outside the outermost horizon in asymptotically flat, anti-de Sitter (AdS), and de Sitter (dS) black hole spacetimes. Our results show that between two neighboring horizons (including the cosmological horizon in the dS case), the topological number is W=1W=-1, indicating the presence of at least one unstable TCO. Outside the outermost horizon, we find $W=0$ for both asymptotically flat and AdS black holes, implying that any TCOs must appear in stable-unstable pairs or be absent. These conclusions are independent of the spin orientation (co-rotating or counter-rotating) of the test particle. The analysis is supported by explicit examples in Schwarzschild, Schwarzschild-AdS, and Schwarzschild-dS spacetimes, confirming the general topological predictions. While the effective potential for spinning particles has been previously studied, the topological approach employed here reveals invariant properties that remain robust even when spin is included, thereby highlighting the fundamental influence of the spacetime structure itself.
I Introduction
Circular orbits in black-hole spacetimes are essential tools for probing strong-field gravity, underpinning key astrophysical phenomena such as accretion disks, black-hole shadows, and gravitational-wave emission from extreme mass-ratio inspirals. Their study provides deep insight into the structure of spacetime and the behavior of matter in extreme environments.
In recent years, topological methods have offered a powerful, model-independent framework for analyzing the existence and stability of circular orbits. These approaches provide a geometric, global perspective on orbital structures that transcends the details of specific metrics or matter content. Cunha, Berti and Herdeiro Cunha:2017qtt pioneered a topological framework to study light rings (LRs) or photon spheres (PSs) of ultra-compact objects, demonstrating that non-degenerate LRs always appear in pairs. They then proceeded to investigate the topological properties of the LRs outside a stationary, axisymmetric black hole Cunha:2020azh , proving that at least one unstable LR exists outside the black hole horizon for each rotation sense. Subsequent work, particularly by Wei et al.  Wei:2020rbh , applied Duan’s topological-current $\phi$-mapping theory Duan:1979ucg ; Duan:1984ws to classify circular orbits in asymptotically flat, AdS and dS black holes, uncovering universal topological invariants associated with photon spheres. Duan’s topological current $\phi$-mapping can also provide new insights into the study of black hole and black hole thermodynamics Wei:2021vdx ; Wei:2022dzw .
Although significant progress has been made in understanding null circular orbits, the case of timelike circular orbits (TCOs) is more intricate, owing to their dependence on orbital parameters such as angular momentum. Earlier studies established that, for neutral particles with fixed angular momentum, TCOs generically arise in stable–unstable pairs Wei:2022mzv ; Ye:2023gmk ; this pairing is topologically enforced and independent of black-hole parameters. Building on this, several further studies have been carried out Shahzad:2024oxb , including the case of charged particles Ye:2024sus ; Song:2025vhw .
In realistic scenarios, however, particles can possess intrinsic spin. The motion of spinning test particles is governed by the Mathisson–Papapetrou–Dixon (MPD) equations Mathisson:1937zz ; Papapetrou:1951pa ; Tulczyjew:1959 ; Taub:1964 ; Madore:1969 ; Dixon:1964 ; Dixon:1965:1 ; Dixon:1970:1 ; Dixon:1970:2 ; Dixon:1973:1 ; Dixon:1974:1 ; Dixon:1979 ; Obukhov:2015eqa ; Semerak:1999qc ; Hackmann:2014tga ; Steinhoff:2009tk , which describe the dynamics of pole–dipole particles in curved spacetime. Spin–curvature coupling induces non-geodesic motion, enriching orbital dynamics and producing spin-dependent shifts in orbital energies and radii. The circular orbits of a spinning test particle in different spacetimes have been studied Jefremov:2015gza ; Toshmatov:2019bda ; Toshmatov:2020wky .
A natural question therefore arises: how does a test particle’s spin affect the topological structure of its circular orbits? Does spin alter the topological invariants that characterise orbital stability? Moreover, are these results universal across different asymptotic behaviours—flat, AdS or dS—and in black hole with multiple horizons? We address these questions by employing a topological approach to study the circular orbits of spinning test particles in static, spherically symmetric spacetimes. Using the Tulczyjew spin-supplementary condition [29], we construct an auxiliary potential and an associated vector field on the equatorial plane. By computing the topological winding number $W$, we classify the existence and stability of TCOs in various regions of black-hole spacetimes.
Our results show that the topological number $W$ is independent of both the magnitude and the orientation of the spin—that is, whether the particle is co-rotating or counter-rotating. Specifically, between two neighbouring horizons we obtain W=1W=-1, indicating the presence of at least one unstable TCO. Outside the outermost horizon of asymptotically flat and AdS spacetimes we find $W=0$, implying that any TCOs must appear in stable–unstable pairs or be absent altogether. These conclusions are corroborated by explicit examples in Schwarzschild, Schwarzschild–AdS and Schwarzschild–dS spacetimes.
This study extends the topological classification of circular orbits to spinning test particles, revealing a robust topological structure that persists even in the presence of spin-induced non-geodesic effects. It underscores the universality of topological methods in gravitational physics and opens new avenues for exploring the dynamics of spinning bodies in more complex spacetimes.
The present paper is organized as follows: In Sec. II, we will give a brief review of the equations of motion of spinning extended test bodies, i.e., MPD equations, and introduce Wei et al.’s topological approach to the study of circular orbits. In Sec. III, we will study the topology of circular orbits for spinning test particles between two neighboring horizons. In Sec. IV, we will study the topology of circular orbits for spinning test particles in asymptotically flat, AdS, and dS black holes. In Sec. V, we summarize and discuss our results. In this work, we use geometrized units, setting $G=c=1$.
II TOPOLOGICAL APPROACH FOR CIRCULAR ORBITS
In this section, we present the general formalism for a spinning test particle within the Mathisson–Papapetrou–Dixon (MPD) approximation, up to the pole-dipole order Mathisson:1937zz ; Papapetrou:1951pa ; Dixon:1964 ; Dixon:1965:1 ; Dixon:1970:1 ; Dixon:1970:2 ; Dixon:1973:1 ; Dixon:1974:1 ; Dixon:1979 and review the topological approach for studying the properties of circular orbits. We consider a spinning particle moving in a static, spherically symmetric spacetime, with the line element:
$ds^{2}=-f(r)dt^{2}+f(r)^{-1}dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2})\;,$ (1)
where $f$ is the function of radial coordinate $r$. The equations of motion for spinning test particles up to the pole-dipole order are given by the MPD equations:
$\dot{P}^{a}=-\frac{1}{2}R^{a}{}_{bcd}u^{b}S^{cd}\;,$ (2)
$\dot{S}^{ab}=2P^{[a}u^{b]}\;.$ (3)
where $D/d\lambda$ is the covariant derivative along the particle’s trajectory with the affine parameter $\lambda$ given by $D/d\lambda\equiv u^{a}\nabla_{a}$ and $R^{a}{}_{bcd}$ is the Riemann tensor. The dynamical 4-momentum and kinematical 4-velocity of the particle are denoted by $p^{a}$ and $u^{a}$, respectively and the anti-symmetric spin tensor is denoted by $S_{ab}$ (with $S_{ab}=-S_{ba}$).
To close the system of Eqs. (2) and (3), a supplementary condition must be imposed. In this work, to restrict the spin tensor to generate rotations only, we use the Tulczyjew spin-supplementary condition Tulczyjew:1959 , i.e.,
$S^{ab}P_{b}=0\;.$ (4)
From Eq.(4), it follows that the canonical momentum and the spin of the particle provide two independent conserved quantities:
$P^{a}P_{a}=-\mathcal{M}^{2}\;,$ (5)
$\frac{1}{2}S^{ab}S_{ab}=S^{2}\;,$ (6)
where $\mathcal{M}$ is the ‘dynamical’, ‘total’ or ‘effective’ rest mass of the body and $S$ is the spin length of the particle. The spin four-vector can be defined as
$S^{a}=\frac{1}{2\mathcal{M}}\epsilon^{ba}{}_{cd}P_{b}S^{cd}\;,$ (7)
where $\epsilon^{ba}{}_{cd}$ is the Levi–Civita tensor. It is easy to find out $S^{a}$ is orthogonal to $P^{a}$, i.e., $S^{a}P_{a}=0$.
The Tulczyjew spin-supplementary condition (4) implies that the components of the 4-velocity $u^{a}$ are determined from the following relation Kuenzle:1972uk
$u^{a}=\frac{m}{\mathcal{M}^{2}}\bigg(P^{a}+\frac{2S^{ab}R_{bcde}P^{c}S^{de}}{4\mathcal{M}^{2}+R_{abcd}S^{ab}S^{cd}}\bigg)\;,$ (8)
where $m$ is a scalar parameter (the ‘kinematical’ or ‘monopole’ rest mass of a particle), and it is given by $u^{a}P_{a}=-m$.
The conserved quantities associated with the spacetime symmetries via the Killing vectors $\xi^{a}$ can be expressed as
$P^{a}\xi_{a}-\frac{1}{2}S^{ab}\nabla_{b}\xi_{a}=P^{a}\xi_{a}-\frac{1}{2}S^{ab}\partial_{b}\xi_{a}=\mathrm{constant}\;.$ (9)
Due to the spherical symmetry of the metric (1), and provided the spin vector is aligned perpendicular to the equatorial plane (parallel to the total angular momentum) Hackmann:2014tga , we can confine the motion to the equatorial plane, i.e., $\theta=\pi/2$. Along the worldline, there are two conserved quantities for the pole-dipole particle: the energy $E$ and the angular momentum $L$. From Eq.(9), the conserved quantities can be expressed as
$-E=P_{t}+\frac{1}{2}f^{\prime}S^{tr}\;,$ (10)
$L=P_{\phi}+rS^{r\phi}\;.$ (11)
where a prime denotes the derivative with respect to radial coordinate $r$. From the Tulczyjew spin supplementary condition (4), and noting that for motion confined to the equatorial plane the components, i.e., $S^{\mu\theta}=0$ Hackmann:2014tga , we obtain:
$S^{t\phi}=-\frac{P_{r}}{P_{\phi}}S^{tr}\;,$ (12)
$S^{r\phi}=\frac{P_{t}}{P_{\phi}}S^{tr}\;.$ (13)
From Eq.(5), we can get
$(P^{r})^{2}=P_{t}^{2}-f\bigg(\frac{P_{\phi}^{2}}{r^{2}}+\mathcal{M}^{2}\bigg)\;.$ (14)
From Eq.(6), and considering Eqs.(12), (13) and (14), one finds Hackmann:2014tga ; Toshmatov:2020wky
$S^{tr}=\frac{sP_{\phi}}{r}\;,$ (15)
where $s=S/\mathcal{M}$ is specific spin parameter. It should be noted that $s$ can have both negative and positive values depending on the direction of spin with respect to the direction of the orbital angular momentum $L$. While $L$ is conserved, $P_{\phi}$ is not necessarily constant, see Eq. (11). The sign of $s$ relative to $L$ defines co/counter-rotation. In the spherically symmetric case, $L$ can always be chosen to be greater than zero, i.e., $L>0$. Then, when $s<0$, the spin and angular momentum are counter-rotating; when $s>0$, the spin and angular momentum are co-rotating. Toshmatov:2020wky . From the conservation of energy (10) and angular momentum (11), we have
$P_{t}=-\frac{2rE+f^{\prime}Ls}{2r-f^{\prime}s^{2}}\;,$ (16)
$P_{\phi}=\frac{2r(L+Es)}{2r-f^{\prime}s^{2}}\;.$ (17)
To ensure the mathematical consistency of the derived expressions (e.g., avoiding divergences in $p_{\phi}$, see Eq.(16)) and to maintain the timelike character of the particle’s four-velocity $u^{a}$, see Eq.(8), (thus preserving causality outside the event horizon), we restrict our analysis to cases where the spin magnitude is small enough such that [^4] This condition, which appears in the denominator of key quantities like Eq.(16), helps prevent the spin-curvature coupling from inducing significant deviations that might violate the timelike character of $P^{a}$ and $u^{a}$. While $P^{a}P_{a}$ is timelike by construction from the Tulczyjew condition ($P^{a}P_{a}=-\mathcal{M}^{2}$), the four-velocity $u^{a}u_{a}$ is determined by a more complex relation involving $P^{a}$ and the spin-curvature coupling term. If the spin becomes too large, the denominator $2r-f^{\prime}s^{2}$ can vanish or become negative, leading to divergences or sign changes in momenta components ($P_{t},P_{\phi}$) and potentially resulting in a spacelike or null , implying superluminal motion. This condition is a sufficient (but not always strictly necessary) constraint to avoid these unphysical scenarios and ensure we work within the regime where the pole-dipole approximation and our analysis are valid. While this holds for the spin magnitudes and regions considered, extreme values of $|s|$ (comparable to the black hole mass $M$) or regions very close to the horizon in near-extremal black holes may violate it. Such regimes require a different treatment, which is beyond the scope of this analysis.:
$2r-f^{\prime}s^{2}>0\;.$ (18)
Putting eq.(16) and (17) into eq.(14), we get the result that
$(P^{r})^{2}\equiv\mathcal{V}=A(E-V_{+})(E-V_{-})\;,$ (19)
where
$A=\frac{4(r^{2}-fs^{2})}{(2r-f^{\prime}s^{2})^{2}}\;,$ (20)
and
$V_{\pm}=\frac{(2f-rf^{\prime})sL}{2(r^{2}-fs^{2})}\pm\frac{2r-f^{\prime}s^{2}}{2(r^{2}-fs^{2})}\sqrt{f[L^{2}+\mathcal{M}^{2}(r^{2}-fs^{2})]}\;.$ (21)
Let $B=r^{2}-fs^{2}$. At the horizon, where $f=0$, we have $B>0$. Outside the horizon, based on the assumption (18), we have $B^{\prime}=2r-f^{\prime}s^{2}>0$. Therefore, $B$ is monotonically increasing outside the horizon, which implies that $B>0$ everywhere outside the horizon. Consequently, $A>0$ everywhere outside the horizon. For $(P^{r})^{2}\geq 0$, the energy must satisfy:
$E\geq V_{+}\;,\quad\mathrm{or}\quad E\leq V_{-}\;,$ (22)
We focus on the case of positive energy, which corresponds to the effective potential $V=V_{+}$. The conditions for circular orbits, $\mathcal{V}=\partial_{r}\mathcal{V}=0$, become:
$E=V\;,\quad\mathrm{and}\quad\partial_{r}V=0\;.$ (23)
Similar to Ref. Wei:2022mzv , $\partial_{r}V$ depends on $r$, $L$, and $s$, A regular potential function, which is independent of the orbital parameters, cannot be introduced here, as discussed in Cunha:2017qtt ; Cunha:2020azh . For simplicity, we do not consider the four-dimensional expressions involving $\theta$. Instead, following the approach of choosing auxiliary functions in Ref. Wei:2021vdx , we introduce a new function
$V_{\theta}=\frac{1}{\sin\theta}V\;.$ (24)
where the factor $\frac{1}{\sin\theta}$ is an auxiliary term. This construction is motivated by the need to define a vector field $\phi^{a}$ (see below) that is regular on the entire $(r,\theta)$ plane, including the poles $\theta=0,\pi$, and whose zeros correspond to circular orbits on the equatorial plane. The factor $1/\sin\theta$ ensures the asymptotic behavior of $\phi^{a}$ at the boundaries $\theta=0,\pi$ is universal and independent of $L$ and $s$, which is crucial for a consistent topological classification.
In order to give a global topology, we require that the values of the angular momentum and spin do not change the asymptotic behavior of $\partial_{r}V_{\theta}$ at the boundary of the $(r,\theta)$ plane. Following the Ref. Wei:2020rbh , we define a vector field $\phi=(\phi^{r},\phi^{\theta})$
$\phi^{r}=\frac{\partial_{r}V_{\theta}}{\sqrt{g_{rr}}},\quad\phi^{\theta}=\frac{\partial_{\theta}V_{\theta}}{\sqrt{g_{\theta\theta}}}\;,$ (25)
in a flat vector space. From Eqs.(21) (24) and (25), we have
$\phi^{r}$ $=\frac{1}{4(r^{2}-s^{2}f)^{2}}\bigg\{2Ls\sqrt{f}[(2f-rf^{\prime})(f^{\prime}s^{2}-2r)+(r^{2}-s^{2}f)(f^{\prime}-rf^{\prime\prime})]$
$-2f[(2r-s^{2}f^{\prime})^{2}+(r^{2}-s^{2}f)(s^{2}f^{\prime\prime}-2)]\sqrt{L^{2}+\mathcal{M}^{2}(r^{2}-s^{2}f)}$
$+\frac{(r^{2}-s^{2}f)(2r-s^{2}f^{\prime})[(L^{2}+r^{2}\mathcal{M}^{2})f^{\prime}+2\mathcal{M}^{2}f(r-s^{2}f^{\prime})]}{\sqrt{L^{2}+\mathcal{M}^{2}(r^{2}-s^{2}f)}}\bigg\}\;,$ (26)
and
$\phi^{\theta}=-\frac{\cos\theta}{r\sin^{2}\theta}\bigg[\frac{(2f-rf^{\prime})sL}{2(r^{2}-fs^{2})}+\frac{2r-f^{\prime}s^{2}}{2(r^{2}-fs^{2})}\sqrt{f(L^{2}+\mathcal{M}^{2}(r^{2}-fs^{2}))}\bigg]\;.$ (27)
It follows that
[…]
$\phi^{r}=\phi\cos\Omega\;,\quad\phi^{\theta}=\phi\sin\Omega\;,\quad\Omega=\arctan(\phi^{\theta}/\phi^{r})\;.$ (29)
For each circular orbit, we assign a topological charge. Following Refs. Duan:1979ucg ; Duan:1984ws ; Wei:2020rbh , the topological current is:
$j^{\mu}=\frac{1}{2\pi}\epsilon^{\mu\nu\rho}\epsilon_{ab}\partial_{\nu}n^{a}\partial_{\rho}n^{b}\;,$ (30)
[…]
$j^{\mu}=\delta^{2}(\phi)J^{\mu}\bigg(\frac{\phi}{x}\bigg)$ (31)
with the Jacobi tensor:
$\epsilon^{ab}J^{\mu}\bigg(\frac{\phi}{x}\bigg)=\epsilon^{\mu\nu\rho}\partial_{\nu}\phi^{a}\partial_{\rho}\phi^{b}\;.$ (32)
Note that $J^{\mu}$ is nonzero only at the zeros of $\phi$. The topological current density is:
$j^{0}=\sum_{i}^{N}=\beta_{i}\eta_{i}\delta^{2}(\vec{x}-\vec{z}_{i})\;,$ (33)
where the Hopf index ($\beta_{i}$) and the Brouwer degree ($\eta_{i}$) of the $i$-th zero point are expressed. The topological number is:
$W=\int_{\Sigma}j^{0}d^{2}x=\sum_{i}^{N}\beta_{i}\eta_{i}=\sum_{i}^{N}w_{i}\;,$ (34)
where $w_{i}$ is the winding number of the $i$-th zero point within region $\Sigma$. Alternatively, $W$ can be computed as:
$W=\frac{1}{2\pi}\oint_{I}d\Omega\;.$ (35)
where $I=\partial\Sigma$ is a counterclockwise closed contour enclosing $\Sigma$. We choose the contour $C=\sum_{i}\cup l_{i}$ as shown in Fig. 1.
To compute $W$, we analyze the behavior of Eqs.(II) and (27) on the boundaries. At $\theta=0$ and $\pi$, we have:
$\phi^{r}_{l_{4}}(\theta\to 0^{+})\sim\frac{1}{\theta},\quad\phi^{\theta}_{l_{4}}(\theta\to 0^{+})\sim-\frac{1}{\theta^{2}}\;,\phi^{r}_{l_{2}}(\theta\to\pi^{-})\sim\frac{1}{\pi-\theta},\quad\phi^{\theta}_{l_{2}}(\theta\to\pi^{-})\sim\frac{1}{(\pi-\theta)^{2}}\;,$ (36)
Thus, the vector $\phi$ points vertically upward at $\theta=0$ and vertically downward at $\theta=\pi$, which is shown in Fig.1. The asymptotic behavior of $\phi^{r}$ and $\phi^{\theta}$ at $l_{1}$ and $l_{3}$ will be studied in the following sections based on our previous work Song:2025vhw .
Figure 1: Representation of the contour $C=\sum_{i}\cup l_{i}$ (which encloses $\Sigma$) on the $(r,\theta)$ plane. The curve $C$ has a positive orientation, marked with the red arrows. $r_{\mathrm{in}}$ and $r_{\mathrm{out}}$ have different meanings in different cases. The black arrows indicate the approximate directions of the vector $\phi$ at the boundaries. At $\theta=0$ and $\pi$, the direction of the vector $\phi$ is vertically upward and downward, respectively.
III Between two neighboring horizons
In this section, we study the topology of circular orbits for spinning test particles between two neighboring horizons.
Assume multiple horizons exist, and consider two neighboring horizons $r_{h_{1}}$ and $r_{h_{2}}$ with $r_{h_{1}}<r_{h_{2}}$. The behavior of $f(r)$ is shown in Fig. 2. We consider only case (a), where $f(r)>0$ between the horizons. In case (b), $f(r)<0$ would make $\phi^{\theta}$ and $\phi^{r}$ undefined, so circular orbits are forbidden. At the horizon, one has $f(r_{h})=0$.
(a).  $f^{\prime}>0$ at $r_{h_{1}}$, $f^{\prime}<0$ at $r_{h_{2}}$ and $f^{\prime\prime}<0$.
(b).  $f^{\prime}<0$ at $r_{h_{1}}$, $f^{\prime}>0$ at $r_{h_{2}}$ and $f^{\prime\prime}>0$.
Figure 2: The behavior of $f(r)$ in the region between two neighboring horizons.
We adopt the contour $C$, where $C=\sum_{i}l_{i}$ where $l_{1}\sim l_{4}$$\{r_{\mathrm{out}}=r_{h_{2}},0\leq\theta\leq\pi\}\cup\{\theta=\pi,r_{h_{1}}\leq r<r_{h_{2}}\}\cup\{r_{\mathrm{in}}=r_{h_{1}},0\leq\theta\leq\pi\}\cup\{\theta=0,r_{h_{1}}\leq r<r_{h_{2}}\}$, as illustrated in Fig. 3.
Figure 3: Representation of the contour $C=\sum_{i}l_{i}$ (which encloses $\Sigma$) on the $(r,\theta)$ plane. The curve $C$ has a positive orientation, marked with the red arrows. The black and blue arrows indicate the approximate directions of the vector $\phi$ at the boundaries.
At $r_{h}$, on the equatorial plane, we have:
$\phi^{r}_{l_{3}}(r\to r_{h})\sim\frac{f^{\prime}(2r-f^{\prime}s^{2})\sqrt{L^{2}+\mathcal{M}^{2}r^{2}}}{4r^{2}}$ (37)
For the case (a) in Fig. 2, at $r_{h_{1}}$, we have $f^{\prime}(r_{h_{1}})>0$. From Eq.(18), we find:
$\phi^{r}_{l_{3}}(r\to r_{h_{1}}^{+})>0\;,$ (38)
Ignoring the specific value of $\phi^{\theta}$, the vector $\phi$ points to the right on $l_{3}$.
At $r_{h_{2}}$, since $f^{\prime}(r_{h_{2}})<0$, we have
$\phi^{r}_{l_{1}}(r\to r_{h_{2}}^{-})<0\;,$ (39)
so $\phi$ points to the left on $l_{1}$.
These results are shown by blue arrows in Fig. 3. Combining the behavior at $\phi$ at $\theta=0$ and $\theta=\pi$, we obtain
$W=\frac{1}{2\pi}\oint d\Omega=\frac{1}{2\pi}\times(-2\pi)=-1\;.$ (40)
This indicates that for fixed $L$ and $s$, there is always at least one unstable TCO between $r_{h_{1}}$ and $r_{h_{2}}$ Cunha:2020azh ; Wei:2022mzv . This result is independent of the particle’s spin and orientation.
IV Outsider the outermost horizon
IV.1 Asymptotically flat black holes
In an asymptotically flat black hole described by (1), the metric function behaves as Wei:2020rbh :
$f\sim 1-\frac{2M}{r}+\mathcal{O}\bigg(\frac{1}{r^{2}}\bigg)\;,r\to\infty\;,$ (41)
where $M$ is the black hole mass. For $r_{h}<r<\infty$, $f(r)>0$ and $f^{\prime}(r_{h})>0$, as shown in Fig. (4).
Figure 4: The behavior of $f(r)$ in an asymptotically flat black hole. At $r_{h}$, one has $f^{\prime}(r)>0$.
We choose the contour $C=\sum_{i}l_{i}$ as $\{r_{\mathrm{out}}=\infty,0\leq\theta\leq\pi\}\cup\{\theta=\pi,r_{h}\leq r<\infty\}\cup\{r_{\mathrm{in}}=r_{h},0\leq\theta\leq\pi\}\cup\{\theta=0,r_{h}\leq r<\infty\}$, where $r_{h}$ is the outermost horizon. This is shown in Fig. 5.
Figure 5: Representation of the contour $C=\sum_{i}l_{i}$ (which encloses $\Sigma$) on the $(r,\theta)$ plane. The curve $C$ has a positive orientation, marked with the red arrows. The black, blue and yellow arrows indicate the approximate directions of the vector $\phi$ at the boundaries.
At $\infty$, on the equatorial plane:
$\phi_{l_{1}}^{r}(r\to\infty)\sim\frac{M\mathcal{M}}{r^{2}}$ (42)
so $\phi$ points to the right on $l_{1}$ (yellow arrow).
At $r_{h}$, from Eq.(37):
$\phi^{r}_{l_{3}}(r\to r_{h}^{+})>0\;,$ (43)
so $\phi$ points to the right on $l_{3}$ (blue arrow). Combining with the vertical directions at $\theta=0$ and $\theta=\pi$, we find:
$W=\frac{1}{2\pi}\oint d\Omega=\frac{1}{2\pi}\times(\pi-\pi)=0\;.$ (44)
This implies that any TCOs with fixed $L$ and $s$ must appear in stable-unstable pairs or be absent. This result is independent of spin.
  • Example: Schwarzschild black hole
For a Schwarzschild black hole:
$f(r)=1-\frac{2M}{r}\;.$ (45)
Set $M=1$ and $\mathcal{M}=0.5$. The horizon is located at $r_{h}=2$. The effective potential becomes
$V(r)=\frac{Ls(r-3M)r^{2}+(r^{3}-Ms^{2})\sqrt{(r-2M)[L^{2}r+\mathcal{M}^{2}(r^{3}+2Ms^{2}-rs^{2})]}}{r^{2}(r^{3}+2Ms^{2}-rs^{2})}\;.$ (46)
  • (1).
    For $L=0.3$ and $1\leq s\leq 1$, the condition $2r-f^{\prime}s^{2}|_{r_{h}}>0$ is satisfied. No TCOs exist (Fig. 6).
    Figure 6: The graph of the potential function (46) for $L=0.3$ and $1\leq s\leq 1$. The red points and the blue boxes represent the maximum values and the minimum values of (46) for a given value of $s$, respectively.
  • (2).
    For $L=1$ and $1\leq s\leq 1$, the condition $2r-f^{\prime}s^{2}|_{r_{h}}>0$ is satisfied. Pairs of stable and unstable TCOs exist (Fig. 7).
    Figure 7: The graph of the potential function (46) for $L=1$ and $1\leq s\leq 1$. The red points and the blue boxes represent the maximum values and the minimum values of (46) for a given value of $s$, respectively.
The numerical evidence from the graphical results for both $L=0.3$ (no TCOs) and $L=1$ (stable-unstable TCO pairs) confirms the topological conclusion $W=0$, which is independent of spin [^5] The angular momentum $L$ must exceed a critical value to provide sufficient centrifugal support for the existence of stable-unstable TCO pairs, otherwise, no TCOs exist..
IV.2 Asymptotically AdS black holes
For an asymptotically AdS black hole Wei:2020rbh :
$f(r)\sim\frac{r^{2}}{l^{2}}+1-\frac{2M}{r}+\mathcal{O}\bigg(\frac{1}{r^{2}}\bigg)\;,$ (47)
where $l$ is the AdS radius. For $r_{h}<r<\infty$, $f(r)>0$ and $f^{\prime}(r_{h})>0$ (Fig. 8).
Figure 8: The behavior of $f(r)$ in an asymptotically AdS black hole. At $r_{h}$, one has $f^{\prime}(r)>0$.
The contour is the same as in the asymptotically flat case (Fig. 9)
Figure 9: Representation of the contour $C=\sum_{i}l_{i}$ (which encloses $\Sigma$) on the $(r,\theta)$ plane. The curve $C$ has a positive orientation, marked with the red arrows. The black, blue and yellow arrows indicate the approximate directions of the vector $\phi$ at the boundaries.
At $\infty$,
$\phi_{l_{1}}^{r}(r\to\infty)\sim\frac{\mathcal{M}\sqrt{l^{2}-s^{2}}}{l^{3}}r>0\quad(\mathrm{if}\;l>|s|)\;,$ (48)
so $\phi$ points to the right (yellow arrow).
At $r_{h}$,
$\phi^{r}_{l_{3}}(r\to r_{h}^{+})>0\;,$ (49)
so $\phi$ points to the right (blue arrow). Thus:
$W=\frac{1}{2\pi}\oint d\Omega=\frac{1}{2\pi}\times(\pi-\pi)=0\;.$ (50)
TCOs must occur in pairs or be absent, independent of spin (as long as $l>|s|$).
  • Example: Schwarzschild-AdS (Sch-AdS) Black Hole
For Sch-AdS:
$f(r)=1-\frac{2M}{r}+\frac{r^{2}}{l^{2}}\;,$ (51)
Set $M=1$, $\mathcal{M}=0.5$ and $l=10$. The horizon is located at $r_{h}\approx 1.9283$. The behavior of $f(r)$ in Sch-AdS black hole is shown in Fig. 10.
Figure 10: The behavior of $f(r)$ in Sch-AdS black hole. At $r_{h}$, one has $f^{\prime}(r)>0$.
The effective potential is:
$V(r)=\frac{(r-3M)l^{2}Lsr+[l^{2}r^{3}-(l^{2}M+r^{3})s^{2}]\sqrt{(1-\frac{2M}{r}+\frac{r^{2}}{l^{2}})\{L^{2}+\mathcal{M}^{2}[r^{2}-(1-\frac{2M}{r}+\frac{r^{2}}{l^{2}})s^{2}]\}}}{l^{2}r^{2}[r^{2}-(1-\frac{2M}{r}+\frac{r^{2}}{l^{2}})s^{2}]}\;.$ (52)
  • (1).
    For $L=0.3$ and $0.25\leq s\leq 0.25$, the conditions $2r-f^{\prime}s^{2}|_{r_{h}}>0$ and $l>|s|$ are satisfied. No TCOs exist (Fig. 11).
    Figure 11: The graph of the potential function (52) for $L=0.3$ and $0.25\leq s\leq 0.25$. The red points and the blue boxes represent the maximum values and the minimum values of (52) for a given value of $s$, respectively.
  • (2).
    For $L=2$ and $1\leq s\leq 1$, the conditions $2r-f^{\prime}s^{2}|_{r_{h}}>0$ and $l>|s|$ are satisfied. Pairs of stable and unstable TCOs exist (Fig. 12).
    Figure 12: The graph of the potential function (52) for $L=2$ and $1\leq s\leq 1$. The red points and the blue boxes represent the maximum values and the minimum values of (52) for a given value of $s$, respectively.
The numerical evidence from the graphical results for both $L=0.3$ (no TCOs) and $L=2$ (stable-unstable TCO pairs) confirms the topological conclusion $W=0$, which is independent of spin.
IV.3 Asymptotically dS black holes
For an asymptotically dS black hole Wei:2020rbh :
$f(r)\sim-\frac{r^{2}}{l^{2}}+1-\frac{2M}{r}+\mathcal{O}\bigg(\frac{1}{r^{2}}\bigg)\;,\quad r\to\infty\;,$ (53)
where $l$ is the curvature radius of dS spacetime. Besides the black hole horizon $r_{h}$, there is a cosmological horizon $r_{c}>r_{h}$, with $f(r_{c})=0$. We consider $r_{h}<r<r_{c}$. The behavior of $f(r)$ is as in Fig. 13 (case (a) of Fig. 2). Thus, from Sec. III, we have:
$W=-1\;,$ (54)
indicating at least one unstable TCO for fixed $L$ and $s$.
Figure 13: The behavior of $f(r)$ in asymptotically dS black hole. One has $f^{\prime}(r)>0$ at $r_{h}$ and $f^{\prime}(r)<0$ at $r_{c}$. Here, $r_{h}$ denotes the outermost black hole horizon, while $r_{c}$ represents the cosmological horizon.
  • Example: Schwarzschild-dS (Sch-dS) Black Hole
For Sch-dS:
$f(r)=1-\frac{2M}{r}-\frac{r^{2}}{l^{2}}\;.$ (55)
Set $M=1$, $\mathcal{M}=0.5$ and $l=10$. The graph of the function $f(r)$ is shown in Fig. 14, with $r_{h}\approx 2.09149$ and $r_{c}\approx 8.78885$.
Figure 14: The behavior of $f(r)$ in Sch-AdS black hole. At $r_{h}$, one has $f^{\prime}(r)>0$.
The effective potential is
$V(r)=\frac{(r-3M)rl^{2}Ls+[r^{3}s^{2}+l^{2}(r^{3}-Ms^{2})]\sqrt{(1-\frac{2M}{r}-\frac{r^{2}}{l^{2}})[L^{2}+\mathcal{M}^{2}(r^{2}-s^{2}+\frac{2Ms^{2}}{r}+\frac{r^{2}s^{2}}{l^{2}})]}}{r^{2}l^{2}[r^{2}-(1-\frac{2M}{r}-\frac{r^{2}}{l^{2}})s^{2}]}\;.$ (56)
For $L=0.3$ and $1\leq s\leq 1$, the condition $2r-f^{\prime}s^{2}|_{r_{h}}>0$ satisfies. One unstable TCO exist (Fig. 11).
Figure 15: The graph of the potential function (56) for $L=0.3$ and $1\leq s\leq 1$. The red points and the blue boxes represent the maximum values and the minimum values of (56) for a given value of $s$, respectively.
The numerical evidence from the graphical results for $L=0.3$ (one unstable TCO) confirms the topological conclusion W=1W=-1, which is independent of spin.
V Conclusions and discussions
In this work, we have applied a topological approach to study the circular orbits of spinning test particles in static, spherically symmetric spacetimes. By constructing an auxiliary potential and a vector field $\phi^{a}$ on the equatorial plane, we computed the topological winding number $W$ to classify the existence and stability of TCOs in various black hole spacetimes. Our main conclusions are:
  • (1).
    Between Two Neighboring Horizons: In regions between two horizons (e.g., inner and outer black hole horizons, or between black hole and cosmological horizons in dS spacetime), we find W=1W=-1, indicating at least one unstable TCO for any fixed $L$ and $s$. This generalizes the result for non-spinning particles to spinning ones.
  • (2).
    Outside the Outermost Horizon: In asymptotically flat and AdS spacetimes, we find $W=0$, meaning TCOs must occur in stable-unstable pairs or be absent. Numerical examples in Schwarzschild and Sch-AdS spacetimes confirm this. While the behavior of effective potentials for spinning particles and the existence of circular orbits have been studied since the pioneering work of Tod, de Felice, and Calvani Tod:1976ud , our topological approach reveals invariant properties ($W=0$) that are robust against the inclusion of spin.
  • (3).
    Independence of Spin Orientation: The topological number $W$ is independent of the spin magnitude $s$ and orientation (co- or counter-rotating). Spin affects the location and energy of orbits but not the global topological structure.
This work demonstrates the power of topological methods in classifying circular orbits, even for spinning particles. The invariance of $W$ under spin perturbations suggests a deep robustness in the orbital structure dictated by the background metric. This finding has important implications for gravitational wave astronomy, particularly for modeling extreme mass-ratio inspirals where the secondary object may possess significant spin.
Future directions include extending this analysis to stationary, axisymmetric spacetimes (e.g., Kerr black holes), where the interplay between black hole spin and particle spin may yield richer topological structures. Additionally, studying different spin-supplementary conditions could provide further insights into the gauge dependence of these results. From an astrophysical perspective, it would be valuable to explore the implications of these topological invariants for the gravitational-wave emission from spinning binaries and the formation of accretion disks around black holes.